OpenStudy (princess_dymie):

I need help with this!!!! (1/2)2-6(2-2/3)

OpenStudy (errrwut):

Follow PEMDAS: 1. Parentheses 2. Exponents 3. Multiplication and Division 4. Addition and Subtraction (1/2)*2-6(2-2/3) Do anything in the parentheses first: (2-2/3) = 4/3 (1/2)*2-6(2-2/3) becomes (1/2)*2-6(4/3) now multiply (1/2) *2 = 1 because the 2s cancel each other out 6*(4/3) = 8 (1/2)*2-6(4/3) becomes 1-8 1-8 = -7